### v1

# Comparison of generalised additive models and neural networks in applications: A systematic review

Abstract
Neural networks have become a popular tool in predictive modelling, more commonly associated with machine learning and artificial intelligence than with statistics. Generalised Additive Models (GAMs) are flexible non-linear statistical models that retain interpretability. Both are state-of-the-art in their own right, with their respective advantages and disadvantages. This paper analyses how these two model classes have performed on real-world tabular data. Following PRISMA guidelines, we conducted a systematic review of papers that performed empirical comparisons of GAMs and neural networks. Eligible papers were identified, yielding 143 papers, with 430 datasets. Key attributes at both paper and dataset levels were extracted and reported. Beyond summarising comparisons, we analyse reported performance metrics using mixed-effects modelling to investigate potential characteristics that can explain and quantify observed differences, including application area, study year, sample size, number of predictors, and neural network complexity. Across datasets, no consistent evidence of superiority was found for either GAMs or neural networks when considering the most frequently reported metrics (RMSE, $R^{2}$, and AUC). Neural networks tended to outperform in larger datasets and in those with more predictors, but this advantage narrowed over time. Conversely, GAMs remained competitive, particularly in smaller data settings, while retaining interpretability. Reporting of dataset characteristics and neural network complexity was incomplete in much of the literature, limiting transparency and reproducibility. This review highlights that GAMs and neural networks should be viewed as complementary approaches rather than competitors. For many tabular applications, the performance trade-off is modest, and interpretability may favour GAMs.
Keywords. Neural networks; generalised additive models; systematic review; mixed-effects model.
1 Introduction
Over the past two decades, rapid advancements in machine learning and artificial intelligence have reshaped the abilities of predictive modelling. Among the most transformative of these developments are neural networks (Rosenblatt,, 1958; Rumelhart et al.,, 1986), which have demonstrated high predictive performance across a wide range of tasks, including image recognition, natural language processing, and forecasting (Abiodun et al.,, 2018; Liu et al.,, 2017). In many cases, the primary objective when employing neural networks is to maximise predictive accuracy, with less emphasis placed on model interpretability. These models are often regarded as “black-box” algorithms: although their interconnected and non-linear structure allows for high flexibility, it also makes for a model that is difficult to interpret. One of the key strengths of neural networks is that, even with a single hidden layer, they have been proven to be universal approximators (Cybenko,, 1989; Hornik et al.,, 1989). This property implies that neural networks can approximate any continuous, single-valued function, and hence could be strong predictors. However, these theories of universality may require the neural network size to grow to infinity. Recent studies have shown that deep neural networks of a fixed size, i.e., neural networks with more than one hidden layer, can also achieve universal approximation, often with a drastically reduced number of parameters (Lin et al.,, 2017; Mhaskar and Poggio,, 2016). This depth-induced parsimony improves computational and statistical efficiency.
The majority of neural network research has been conducted outside of the field of statistics, resulting in a limited incorporation of statistically-based methodologies and a disparity in the terminology used (Hooker and Mentch,, 2021). Neural networks, particularly multilayer perceptrons, have emerged as strong alternatives to traditional statistical models, albeit they are often viewed as quite separate entities. However, a number of researchers have illustrated the similarities between neural networks and statistical models. Sarle, (1994) conducted a comprehensive comparison of various neural network models with traditional statistical approaches, showing that multilayer perceptrons can be viewed as generalisations of non-linear regression and discriminant models. Warner and Misra, (1996) highlight the parallels between neural networks and statistical regression models in terms of notation and implementation, showing that neural networks act as a type of nonparametric regression model. Cheng and Titterington, (1994) inform statistical readers about neural networks, highlighting similarities with statistical methodology. More recently, Ha and Burke, (2024) utilise deep neural networks within both a generalised non-linear modelling framework (extending classical generalised linear models) and a non-linear survival model (extending the Cox model). Synthesizing these views, it becomes evident that neural networks are a generalisation of statistical approaches whereby they automatically learn complex non-linear effects and interactions within their hidden layers; this is in contrast to the manual specification of interactions required in traditional statistical modelling, which frequently utilises simpler linear effects for practicality.
Given previous work comparing the mathematical structure of statistical models and neural networks, the question then arises as to how these methods compare with each other in practical applications. Indeed, several authors have made strides at answering this over the years. Sargent, (2001) conducted a review and analysis of the literature comparing neural networks with linear, logistic, and Cox regression in the context of medical research. Karlaftis and Vlahogianni, (2011) provided a survey of applications in the transportation area, comparing neural networks with linear or logistic regression predominantly. Shin et al., (2021) carried out a systematic literature review of machine learning techniques (random forests, decision and regression trees, support vector machines, and neural networks) versus conventional statistical models (logistic, Cox, and Poisson regression) for predicting heart failure readmission and mortality. Cho et al., (2021) carried out a systematic review of the same aforementioned machine learning methods and conventional statistical models for the prediction of myocardial infarction readmission and mortality. Sun et al., (2022) conducted a systematic review and meta-analysis of machine learning techniques, including random forests, support vector machines, neural networks, and decision trees, against statistical methods, linear and Cox regression, for predicting heart failure events. More generally, Paliwal and Kumar, (2009) carried out a review of comparative studies across various application domains, evaluating multilayered feedforward neural networks against linear regression, logistic regression, and discriminant analysis. Their review critically assessed the literature based on several key criteria, including sample size, adherence to the assumptions of the techniques, the method by which results are validated, the metrics used for comparison, and whether statistically significant differences were observed in the outcomes.
Despite the interest in comparing neural networks and statistical models, prior reviews have predominantly narrowed their view to particular subject areas, such as transportation and heart health, as mentioned above. Additionally, these reviews have focused on quite simple statistical models, such as linear regression, which inherently lack the capabilities to model non-linear effects that may be underlying in the data. Therefore, such comparisons are not fair representations of what is possible in more advanced non-linear statistical models. Specifically, Generalised Additive Models (GAMs) (Hastie and Tibshirani,, 1986) are the most popular extension to linear regression, replacing linear terms with arbitrary non-linear functions. The strength of GAMs is their ability to deal with non-linear relationships between predictor and response variables, while maintaining interpretability through their additive structure.
To date, there has been no systematic review of the literature comparing GAMs with neural networks in a wide range of application areas, and, therefore, our current paper closes this gap. The purpose of the systematic review is to be a source of information on the types of studies that have carried out comparisons of these two model classes and the reported findings within these studies. To move beyond just a summary of these papers, we have further extracted and analysed the values of the performance metrics provided within the papers, along with characteristics that can potentially explain observed differences (e.g., application area, year of study, sample size, number of predictor variables, and neural network complexity).
The remainder of the paper is structured as follows. In Section 2, we provide a theoretical background on GAMs and neural networks. Section 3 outlines the research questions guiding this systematic review. The methodology for conducting the review is detailed in Section 4. The study selection process for included papers and a quality assessment of the final sample is provided in Section 5. Then Section 6 summarises the data attributes extracted from the reviewed literature. Analysis and modelling of the performance metrics used to compare GAMs and neural networks is provided in Section 7. Finally, conclusions are presented in Section 8.
2 Theoretical background
Before providing the details of our systematic review, we first provide some brief technical details on the models under consideration. Traditional statistical models and neural networks share the fundamental objective of modelling the relationship between a set of covariates and a response variable. Let $Y\in\mathbb{R}$ denote the response variable and $X=(X_{1},X_{2},\dots,X_{k})^{\top}\in\mathbb{R}^{k}$ the vector of covariates, where we restrict attention to tabular data. The general regression framework can be written as
$\mathbb{E}[Y\mid X=\mathbf{x}]=f(\mathbf{x}),$ (2.1)
where $\mathbb{E}[\cdot]$ is the expectation operator, and $f(\mathbf{x})$ is a function of the observed covariates $\mathbf{x}=(x_{1},x_{2},\dots,x_{k})^{\top}$, encapsulating how the mean of $Y$ depends on $X$. The choice of this function defines the type of regression model employed and may take several well-known forms, each corresponding to a different class of models:
[…]
where GAMs utilise smooth functions $g_{j}(x_{j})$ to model non-linear relationships and
$\textbf{NN}(x_{1},x_{2},\dots,x_{k})$ is a fully connected neural network (NN).
2.1 Generalised additive models (GAMs)
GAMs were first introduced by Hastie and Tibshirani, (1986, 1990), extending linear regression by allowing non-linear relationships between the predictor variables $\mathbf{x}$, and the response $Y$. As seen in Equation 2.2, smooth functions $g(\cdot)$ replace the constant coefficients of linear regression, which increases flexibility. Smooth functions can be approximated as a linear combination of a set of known basis functions,
$g_{j}(x_{j})=\sum_{p=1}^{P}c_{p}b_{p}(x_{j}),$ (2.3)
where $P$ is the number of basis functions, controlling the maximum complexity of the smooth function, $c_{p}$ are constant coefficients to be estimated, and $b_{p}(\cdot)$ are the basis functions. Common basis functions include cubic regression splines (Durrleman and Simon,, 1989), thin-plate regression splines (Wood,, 2003), B-splines (De Boor and De Boor,, 1978), and penalised B-splines (“P-splines”) (Eilers and Marx,, 1996). The estimation of these spline functions, historically achieved via the backfitting algorithm (Hastie and Tibshirani,, 1986), can now be conducted through penalised likelihood maximisation (Wood,, 2011). This approach integrates a roughness penalty directly into the fitting criterion to automatically prevent overfitting. The theoretical and practical development of GAMs has been advanced substantially by Wood’s mgcv (Mixed GAM Computation Vehicle) (Wood,, 2017, 2012) package in the R programming language (R Core Team,, 2024).
A key advantage of GAMs is that their additive nature retains interpretability. By plotting respective smooth functions, GAMs allow the visualisation of each covariate effect. While GAMs have increased flexibility over linear regression, their additive nature, which aids in the interpretation, also limits flexibility. Complex high-degree interactions can be difficult to estimate using splines and are often not specified in practice as a result.
2.2 Neural networks
Neural networks are a class of models inspired by the structure of the human brain, where information flows through the interaction of neurons (Goodfellow et al.,, 2016). Neural networks offer a flexible approach to modelling $f(\mathbf{x})$ by approximating complex, non-linear relationships through layers of interconnected nodes. A standard architecture is the feedforward multilayer perceptron (MLP) neural network. This architecture passes the data through a series of weighted transformations and non-linear activation functions. The default structure of a neural network is fully connected, essentially specifying all possible interactions (assuming a sufficiently complex neural network). An MLP with $L$ layers can be described by the equation:
$f(\mathbf{x})=h_{L}(\mathbf{b}_{L}+\mathbf{W}_{L}h_{L-1}(\dots\mathbf{b}_{2}+\mathbf{W}_{2}h_{1}(\mathbf{b}_{1}+\mathbf{W}_{1}\mathbf{x}))),$ (2.4)
where $L$ is the number of layers, $\mathbf{b}_{\ell}$ is a bias vector, $\mathbf{W}_{\ell}$ is a weight matrix, and $h_{\ell}(\cdot)$ is a non-linear activation function ($\ell=1,\ldots,L$). Figure 1 illustrates the architecture of a feedforward multilayer perceptron neural network. Neural networks are typically optimised via stochastic gradient descent and its variants (Bishop,, 1995; Bottou,, 2010), which iteratively update parameters to minimise a loss function using small, random subsets of data (batches).
The structure of neural networks, while highly flexible, proves difficult to interpret, leading to the label “black-box” models, as understanding the effect of each individual predictor variable and their interactions is often lost in this complexity. The flexibility of these models makes them strong predictors. However, determining key architectural parameters such as the number of hidden layers, the number of nodes per layer, and other hyperparameters is not straightforward. Identifying the optimal configuration typically involves an extensive and computationally intensive search process.
[…]
Figure 1: Feedforward multilayer perceptron neural network.
Table 1: Research questions and motivations.
RQ Research questions Motivation
RQ1 What are the key characteristics of papers carrying out comparisons between GAMs and NNs? Identify domain areas, journals, performance metrics, model structures, and dataset attributes. Determine standard evaluation practices, ensure comparability across studies, and highlight potential research gaps.
RQ2 Which model type performs better? Summarise author-reported conclusion on model performance and advance on this through independent analysis of performance metrics.
RQ3 What are the potential drivers for the difference in performance of GAMs and NNs? Identify factors such as sample size, number of predictors, and dataset characteristics that influence model performance and examine how the relative performance of these models has evolved over time.
3 Research questions
Prior to carrying out a systematic review of comparisons of GAMs and neural networks, we defined key research questions to structure this review and facilitate a targeted comparison between the two model classes. Table 1 highlights our three central research questions along with the motivation behind each.
4 Methods
This review was reported following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines (Page et al.,, 2021).
4.1 Eligibility criteria
To be eligible for inclusion, papers had to be from peer-reviewed journals or conference proceedings, written in English, and have full-text availability. Additionally, papers were required to include original experimental analysis of both GAMs and neural networks using the same dataset. Papers without real-data analysis, such as review, discussion, and theoretical papers, were excluded. Papers in which the values of performance metrics were unattainable or provided in-sample results only were excluded. Duplicate papers and grey literature, including preprints and dissertations, were also excluded.
4.2 Information sources
The search was conducted using Scopus, a large multidisciplinary database of peer-reviewed literature, including journal articles, books, and conference proceedings. The search was limited to titles and abstracts, ensuring that the identified papers were directly relevant to the comparison of GAMs and neural networks. The final search was completed on November 20, 2024, with no restrictions on the earliest publication date, allowing for the inclusion of all relevant literature regardless of publication year.
4.3 Search strategy
To ensure a comprehensive and reproducible search, we formulated a search strategy incorporating synonyms for both GAMs and neural networks. Boolean operators were applied, using ‘OR’ to incorporate variations of each term and ‘AND’ to ensure that both model types were jointly present in the retrieved studies.
The search string used was as follows: ( GAM OR “Generalised additive model” ) AND (“Neural network” OR “artificial-NN” OR “Artificial Neuron Network” OR “Multilayer perceptron” OR MLP))
The systematic search resulted in 445 documents that satisfied the search criteria.
4.4 Selection process
To determine whether a document would be included in the final systematic review, each study was evaluated against the set of predefined eligibility criteria (Section 4.1) to ensure relevance. In order to assess whether a document satisfied these conditions, the entire document was reviewed. The first author conducted the screening of each paper to determine its eligibility for inclusion. In cases where inclusion was uncertain, the paper was discussed with the second and third authors, and a decision was made. This resulted in a total of 143 papers, from which comparisons on 430 datasets were included in our final systematic review (as many papers conduct comparisons on more than one dataset).
4.5 Data collection process
The information collected from the papers in our systematic review is summarised in Table 2. The corresponding source column indicates whether the information was provided by the Scopus search engine, Scimago Journal Rank (SJR), or manually extracted from within the paper itself. Domain area was extracted from Scopus, and then manually combined into higher-level categories by the authors. When studies reported results from multiple neural networks or GAMs, for example, with varying hyperparameters or architectures within the same dataset, information was collected from the model with the best performance. Note, from Table 2 (“Level” column), that there are paper-level attributes (e.g., title and authors), and dataset-level attributes for each of the datasets analysed within a given paper (e.g., author-reported outcome and data type).
Table 2: Data attributes for which information was sought from papers.
Attribute Description Source Level
Title Paper title Scopus Paper
Authors Authors list Scopus Paper
Year Publication year Scopus Paper
Source/journal title Publication source Scopus Paper
Source/journal rank SJR ranking SJR Paper
Domain Application area Scopus/Manual Paper
Problem type Regression or classification task Manual Dataset
Data type Real or simulated data Manual Dataset
Validation Method used, e.g., train-test Manual Dataset
Sample size Number of observations Manual Dataset
Predictors Number of predictor variables Manual Dataset
Neural network type Neural network type, e.g., MLP Manual Dataset
Neural network complexity Layers and nodes count Manual Dataset
Author-reported outcome Author’s conclusion on best model Manual Dataset
Performance metrics Performance metrics used and their values Manual Dataset
5 Study selection
The flow diagram in Figure 2 summarises the study selection process. The search identified 445 papers, including one duplicate, which was removed. Full-text of the remaining 444 papers was screened, of which 143 were included in the review.
Papers identified through database searching ($n=445$) Papers after duplicates removed ($n=444$) Full-text papers assessed for eligibility ($n=444$) Papers included in synthesis ($n=143$) Datasets included in synthesis ($n=429$) Identificationg Screening Includedg Duplicates removed ($n=1$) Full-text papers excluded ($n=301$): Papers not in English ($n=16$) Papers without empirical analysis between GAMs and NNs on same data ($n=235$) Papers lacking numeric values of performance metrics and/or author-reported statement on performance ($n=17$) Papers where results are in-sample ($n=9$) Papers where results are from simulated datasets only ($n=12$) Papers where full text was unavailable ($n=12$)
Figure 2: Flow diagram summarising the results of the paper identification, screening, and inclusion.
5.1 Quality assessment
Among the 15 attributes listed in Table 2, sample size, number of predictors, and neural network complexity were the three that were not always reported in the papers, leading to missingness. The frequency distribution of combinations of these variables and the quantity of datasets with each combination can be seen in Table 3. Only 21% of datasets report all three variables. However, 79% report at least two, and 95% report at least one. Unreported neural network complexity accounts for the majority of missingness, which is reported in only 35% of datasets.
The level of missingness in the literature raises concerns about the level of transparency of reporting in the current literature comparing GAMs to neural networks. Key methodological details, such as dataset characteristics and model complexity, are essential for evaluating the observed performance differences and for the reproducibility of the analyses carried out within these papers. The lack of standardised reporting also suggests a need for clearer guidelines and policies to promote complete reporting of experimental details in publications.
Table 3: Cross-tabulation of availability in collected attributes.
Sample size Predictors Complexity Datasets %
89 21
189 44
39 9
23 5
31 7
39 9
0 0
20 5
332 356 151
77% 83% 35 %
6 Study characteristics
Table 4 lists all 143 papers included in this systematic review, grouped by domain area.
6.1 Paper-level characteristics
Year, publication source, publication rank, and domain area are four key paper-level characteristics that have been gathered from this review. Figure 3 plots the number of papers by domain area through the years of publication. There is a strong increasing trend in papers carrying out comparisons between GAMs and neural networks. The clear surge of activity in recent years emphasises the importance and timeliness of investigating this topic.
There are 121 distinct journals and conference proceedings from which the 143 papers originate. Table 5 summarises the most frequently occurring publication sources, revealing that only six journals appear at least three times. This indicates a broad range of research across multiple disciplines, reflecting the interdisciplinary nature of research comparing GAMs and neural networks. To assess the academic impact of these publications, we analysed their rankings according to the Scimago Journal & Country Rank (SJR) 2024. The distribution is as follows: 69% of publications appear in Q1 (top quartile) journals, 13% in Q2, 8% in Q3, 3% in Q4, and 8% are currently unranked. The majority of comparative studies between GAMs and neural networks are published in Q1 and Q2 sources, reinforcing the significance of this research area.
Table 4: Papers included in the systematic review.
Domain area Papers Datasets
Environmental & Agricultural Sciences 66 199
Aguilos et al., (2024), Beaugrand, (2024), Bonsoms and Ninyerola, (2024), Fan et al., (2024), Fu et al., (2024), Gazis et al., (2024), Jung et al., (2024), Lei et al., (2024), Mouta et al., (2024), Senaratna et al., (2024), Yang et al., (2024), Ghorbani et al., (2024), Adamova and Ukrainskiy, (2023), Boudreault et al., (2023), DelSanto et al., (2023), Li et al., (2023), Yan et al., (2023), Benemann et al., (2022), Shafizadeh-Moghadam et al., (2022), Yadav et al., (2022), Evans and Shen, (2021), Ma et al., (2021), Martín et al., (2021), Mondal and Bhat, (2021), Phelps and Woolford, (2021), Tian et al., (2021), Zhang et al., (2021), Dodangeh et al., (2020), Dorich et al., (2020), Mandal et al., (2020), Yadav et al., (2020), Mugo and Saitoh, (2020), Nguyen et al., (2019), Liu et al., (2018), Pourghasemi and Rahmati, (2018), Corona-Núñez et al., (2017), Gao et al., (2017), Pôças et al., (2017), Ouarda et al., (2016), Shortridge et al., (2016), Vogel and Brown, (2016), Durocher et al., (2015), Li et al., (2015), Lin et al., (2015), Xu et al., (2014), Gutiérrez-Estrada et al., (2012), Queirolo et al., (2012), Haywood and Stone, (2011), Palialexis et al., (2011), Aertsen et al., (2010), Gutiérrez-Estrada and Bilton, (2010), Ogawa-Onishi et al., (2010), Tisseuil et al., (2010), Virkkala et al., (2010), Gutiérrez-Estrada et al., (2009), Lee et al., (2009), López-Darias et al., (2008), Araújo et al., (2005), Luoto and Hjort, (2005), Megrey et al., (2005), Wang et al., (2005), Thuiller, (2003), Brosse and Lek, (2002), Brosse and Lek, (2000), Brosse et al., (1999), Shatar and Mcbratney, (1999)
Engineering, Mathematics, & Computer Science 35 141
Chen et al., (2024), Kang and Choi, (2024), Nascimento et al., (2024), Olca and Yücel, (2024), Pachauri, (2024), Shojaeian et al., (2024), Berrisch et al., (2023), Koh et al., (2023), Ledmaoui et al., (2023), Montero et al., (2023), Razaq and Shukur, (2023), Zhang and Yang, (2023), Zhang et al., (2023), Duan et al., (2022), Mestanza et al., (2022), Pachauri and Ahn, (2022), Dudek, (2021), Fang et al., (2021), Kamis et al., (2021), Matsumoto and Yamada, (2021), Yang et al., (2021), Daniel et al., (2020), Park et al., (2020), Ilseven and Gol, (2019), Tsang et al., (2018), Ilseven and Gol, (2017), Martín et al., (2016), Francisco-Fernández et al., (2015), Yang et al., (2015), Fukuda et al., (2013), Aertsen et al., (2011), Berg, (2007), Schlink et al., (2006), Morlini and Orlandini, (2001), Faraway and Chatfield, (1998)
Health & Biomedical Sciences 12 25
Francisco et al., (2024), Esmaeili et al., (2023), Xu et al., (2023), Moslehi et al., (2022), Shoko et al., (2022), Taylor et al., (2019), Allard et al., (2017), Cengiz and Terzi, (2012), Gregori et al., (2011), Gregori et al., (2009), Giachino et al., (2007), Goldfarb-Rumyantzev and Pappas, (2002)
Other 30 65
Pasha and Reddy, (2024), Roldán et al., (2024), Szeląg et al., (2024), Amara-Ouali et al., (2023), Balan et al., (2023), Bolleddula et al., (2023), Cai et al., (2023), Krämer et al., (2023), Mathur and Mathur, (2023), Sanii and Chan, (2022), Zhou and Zhang, (2022), Dwivedi et al., (2021), Fotiadis et al., (2021), Frasier et al., (2021), Li et al., (2021), Rengarajan et al., (2021), Zhu et al., (2021), Anglart et al., (2020), Zhu et al., (2020), Zihan et al., (2020), Wheeler et al., (2019), Baquero et al., (2018), Luan et al., (2018), Davoudi et al., (2017), Nyhan et al., (2014), Liu and Tesfamariam, (2012), Ryser and Denzler, (2009), Marmion et al., (2008), Heikkinen et al., (2007), Lee et al., (2002)
Figure 3: Number of papers plotted by year of publication, coloured by domain area.
Table 5: Most frequent sources of the papers in this systematic review.
Journal title No. papers SJR best quartile
Ecological Modelling 3 Q1
Science of the Total Environment 3 Q1
Fisheries Research 3 Q1
Environmental Modelling and Software 3 Q1
PLoS ONE 3 Q1
Remote Sensing 3 Q1
High-level domain areas were assigned to each paper based on the areas assigned by Scimago to the source/journal in which the paper appeared. The following four domain areas were used: Environmental and Agricultural Sciences, Engineering, Mathematics, and Computer Science, Health and Biomedical Sciences, and Other. Table 6 summarises domain areas by paper, dataset, and the average number of datasets per paper.
Table 6: Domain areas.
Domain area Papers Datasets Avg. # datasets
Env. & Agri. Sci. (EAS) 66a(46%) 199a(46%) 3
Eng., Math. & Comp. Sci. (EMCS) 35a(24%) 141a(33%) 4
Health & Biomed. Sci. (HBS) 12a(08%) 25a(06%) 2
Other (OTH) 30a(21%) 65a(15%) 2
Total 143 430 3
Abbreviations: EAS = Environmental & Agricultural Sciences; EMCS = Engineering, Mathematics & Computer Science; HBS = Health & Biomedical Sciences; OTH = Other.
6.2 Dataset-level characteristics
Whether the dataset is a regression or classification task, sample size, number of predictors, type of neural network, neural network complexity, author-reported outcome, and performance metrics used are key dataset-level characteristics gathered from papers included in this review. Figure 4 displays a summary of these characteristics.
Regression tasks accounted for 68% of the datasets, while 32% were classification. Sample size, number of predictors, and neural network complexity are three key numerical variables extracted at the dataset level. Descriptive statistics for each are presented in Table 7. These statistics provide insight into the range of the datasets included in the review, as well as the variability in model complexity. As shown, sample sizes range from 35 to over 1.2 million observations, with a median of 1,400. The number of predictors varies widely, from as few as 2 to as many as 3,000, though the median and mean remain low at 7 and 22, respectively. Neural network complexity, defined as the number of parameters, has a wide range from 2 to 120,000 parameters.
Table 7: Summary statistics of dataset-level numerical variables.
Characteristic Min. Median Mean Max. Reported
Sample size 35 1400 15,000 1,200,000 77%
Predictors 2 7 22 3,000 83%
Complexity 2 21 6600 120,000 35%
  • Note: numbers have been displayed to two significant figures.
Figure 4: Proportion of datasets by key characteristics. Reg = regression; Class = classification. Category labels (Small, Medium, Large, Unknown) refer to relative groupings based on sample size, number of predictors, and model complexity.
Feedforward Multilayer perceptron (MLP) neural networks were the most common type, accounting for 89% of datasets. All other neural networks were variations of this, and any one type accounted for less than $5\%$ of the datasets. Neural network complexity was calculated as the total number of parameters, excluding those in the input layer. This exclusion was necessary because input nodes correspond to the number of predictors, which is already captured by the fact that we have collected the number of predictor variables as an attribute. As mentioned, neural network complexity is among the data items frequently unreported. Only 35% of datasets provided sufficient detail to compute the full parameter count. However, 64% reported the number of hidden layers, which includes instances where layers were discussed but not the number of nodes in each needed to look at full complexity. Depth of a neural network via the number of hidden layers is, in its own right, a feature of interest. Among the datasets that did report the number of hidden layers, 70% used a single hidden layer, 22% had two hidden layers, and 7% used between three and five. This distribution indicates that the majority of neural networks used in comparisons with GAMs were shallow architectures, typically with one or two hidden layers.
A wide range of over 80 performance metrics were used across the reviewed datasets. These included well-known metrics such as Root Mean Square Error (RMSE) and the coefficient of determination ($R^{2}$), as well as domain-specific or author-defined measures particular to specific applications. Table 8 displays the ten most frequently used metrics, ranked by the number of datasets in which they appeared. RMSE was the most commonly reported, followed by $R^{2}$, and the Area Under the Receiver Operating Characteristic (AUC) curve.
Table 8: Frequently used performance evaluation measures.
Performance metric Datasets
RMSE 229 (53%)
$R^{2}$ 141 (33%)
AUC 111 (26%)
Mean Absolute Error 086 (20%)
Cohen’s Kappa 040 (09%)
Bias 029 (07%)
True Skill Statistic 022 (05%)
% Accuracy 018 (04%)
$R^{2}$ adjusted 017 (04%)
Relative RMSE 017 (04%)
Looking at author-reported outcomes, neural networks were declared better than GAMs in 27% of datasets. This is more than the 19% of datasets in which GAMs were reported as better than neural networks. In 9% of datasets, both model types were found equivalent. However, author-reported outcomes provide an incomplete and often inconsistent comparison between GAMs and neural networks. Differences can arise in how outcomes are classified from author to author. Additionally, 46% of author outcomes are undefined and hence recorded as “unclear”, often due to a focus on an alternative model in that application. A way to address these limitations is to directly analyse the performance metric values reported for both GAMs and neural networks in each instance. This allows for a more objective comparison and quantifies the magnitude of performance differences, which is not possible when looking at the qualitative author outcomes alone.
7 Performance metric analysis
To quantitatively synthesize the comparative performance of GAMs and neural networks across the reviewed literature, we create a statistical model for the reported performance metric values directly. This analysis examines how the relative performance, measured by the most frequently used metrics, is associated with key study characteristics such as sample size, domain area, and model complexity. Given the hierarchical structure of the data we have collected, where datasets are nested within papers, a mixed-effects modelling approach was employed to account for intra-paper dependence.
7.1 Methods
Mixed-effects models extend linear regression by enabling the analysis of multiple sources of variation in the data, rather than assuming a single source. These models allow the response variable to be modelled through both fixed and random effects. The fixed effects estimate the average response across the entire population, similar to traditional regression, while the random effects (e.g., the influence of individual papers) capture variability across papers and allow for the estimation of paper-specific responses. In this analysis, many papers contain multiple datasets. This introduces a potential dependency (i.e., correlation) in the results, which are likely influenced by the same group of authors applying similar modelling strategies to the datasets they consider. Furthermore, the datasets themselves may be similar, coming from the same (potentially niche) application area or research goal, reducing their independence. Hence, results derived from datasets within the same paper are typically correlated, and this intra-paper dependence must be considered in the analysis. Mixed-effects models capture the correlation between observations in the same group, i.e., datasets within the same paper, known as intraclass correlation (ICC):
$\text{ICC}=\frac{\sigma^{2}_{paper}}{\sigma^{2}_{paper}+\sigma^{2}_{residual}},$ (7.5)
where $\sigma^{2}_{paper}$ is the variation across papers and $\sigma^{2}_{residual}$ is the variation within each paper after accounting for fixed effects and the group intercept. ICC can be interpreted as the correlation between results in datasets from the same paper and also as the proportion of total variance that can be explained by the paper effect. A large ICC indicates the appropriateness of using mixed-effects models compared to standard regression, as the paper effect is strong.
Linear mixed-effects models were fitted using the lme4 package (Bates et al.,, 2015) in R programming language (R Core Team,, 2024). We included a paper-level random intercept effect to account for correlated results within a given paper, along with the following fixed effects: publication year, domain area, sample size, number of predictors, and neural network complexity. Pairwise interactions with year of publication were also included to account for trends over time.
7.1.1 Response and predictor variables
As shown in Table 8, RMSE, $R^{2}$, and AUC are, by far, the most commonly-used performance metrics, accounting for 53%, 33%, and 26% of datasets respectively and collectively covering 89% of all datasets. RMSE and $R^{2}$ are common performance metrics used in regression, accounting for 76% and 48% of regression datasets respectively, and 91% collectively. AUC, a standard metric for classification models, was used in 82% of our classification datasets. These three metrics were selected for our performance metric analysis due to their prevalence and large coverage of datasets.
Figure 5: Log performance ratios plotted by year, where greater than zero signifies neural networks are outperforming GAMs and below zero indicates GAM superiority, as shown by arrows.
The log-ratio of each metric is computed as:
$\log(\mathrm{metric}_{\mathrm{ratio}})=\begin{cases}\log(\mathrm{RMSE}_{\mathrm{GAM}}/\mathrm{RMSE}_{\mathrm{NN}})&\text{for RMSE}\\ \log((1-R^{2}_{\mathrm{GAM}})/(1-R^{2}_{\mathrm{NN}}))&\text{for }R^{2}\\ \log((1-\mathrm{AUC}_{\mathrm{GAM}})/(1-\mathrm{AUC}_{\mathrm{NN}}))&\text{for AUC},\end{cases}$ (7.6)
where positive values indicate superior neural network performance for that metric, negative values signify GAM performance was better, and zero indicates model equivalence. For AUC and $R^{2}$, they are first converted to losses, in which 0 indicates optimal performance, by subtracting them from 1. This allows for all three to represent loss ratios. The ratio of performance metrics provides a relative performance measure that is invariant to scale. This is necessary here, as we are comparing results across many different datasets with substantial variation in absolute performance; different datasets lie on different scales, with differing levels of achievable performance given the inherent noise/uncertainty in the problem at hand. Moreover, we have taken the logarithm to place the response on the real line (so that it is suitable for linear modelling) and to address some skewness. Figure 5 displays the log performance ratios for datasets using RMSE, AUC, and $R^{2}$; points are coloured based on the performance metric used. The horizontal line at zero indicates where GAMs and neural networks are equivalent; points above the line correspond to datasets where neural networks outperform GAMs, while points below are where GAMs achieve superior performance. Overall, we can see that there is a high level of variability in results, with no firm evidence in favour of neural networks or additive models (as the points are generally centred around zero), and no strong trends over time. However, this high-level overview of results does not account for predictor variables (other than the year) or the random effects of papers.
The following predictor variables were considered in the meta-analysis as fixed effects: year of publication (Year), domain area (Area), sample size (Sample_Size), number of predictors (No_Predictors), and neural network complexity (Complexity). Paper ID (Paper_ID) is used as a random effect to account for the dependence between datasets coming from the same papers. To retain as much information as possible, numerical predictors with missing values were converted to categorical variables. Tertiles of the numerical variables were created along with a category “Unknown”, resulting in four levels. This approach enabled the inclusion of incomplete records and allowed assessment of whether missingness was associated with performance outcomes.
Year (numerical, complete):
Values were centred by subtracting the middle year, 2011, from each observation (range: 1998–2024).
Area (categorical, complete):
Classification based on the Scopus categories further grouped into four overarching areas.
Sample_Size (numerical, incomplete):
The number of observations in each dataset. Categorical levels: Small $\leq 350$, Medium (350 - 2,800), Large $\geq 2,800$, and Unknown.
No_Predictors (numerical, incomplete):
The number of predictors in each dataset. Categorical levels: Small $\leq 5$, Medium (5 - 10), Large $\geq 10$, and Unknown.
Complexity (numerical, incomplete):
Number of parameters in the neural network, excluding the input layer. Categorical levels: Small $\leq 6$, Medium (6 - 150), Large $\geq 150$, and Unknown.
Interactions of the predictor variables with year were explored to assess whether or not the effects of these variables change through time. For all models, the category “small” is used as the reference level and 2011 (transformed to be zero) as the reference year.
7.1.2 RMSE metric analysis
The RMSE analysis included 229 dataset observations from 80 papers. After removing two outliers, 227 dataset observations, from 78 papers remained. Including covariates in the analysis allows us to explain more of the variation in RMSE ratios across datasets, and to investigate how aspects of the datasets and models influence the relative performance of neural networks and GAMs. The RMSE model estimates and standard errors (SE) are presented in Table 9. (Models for AUC and $R^{2}$ are also shown, but discussed in Section 7.1.3 and Section 7.1.4 respectively.) The model included all fixed-effects predictors and their pairwise interactions with time (Year) as follows:
$\log(\mathrm{RMSE}_{\mathrm{ratio}})\sim\beta_{0}+\texttt{Year}+\texttt{Area}+\dots+(\texttt{Year}\times\texttt{Complexity})+(1|\texttt{Paper\_ID}),$ (7.7)
where $(1|\texttt{Paper\_ID})$ represents a random intercept capturing variation among datasets. Note that modelling the ratio in this manner is a special case of a slightly more general model discussed in Appendix A.
The intercept for this model is positive, however, not statistically significantly different from zero. This indicates that, at the reference levels for covariates, there is no evidence of a distinction between the performance of GAMs and neural networks. Neural network complexity and its interaction with time significantly affect the log RMSE ratio. From the model output, medium and large complexity effects are both negative and significant, indicating that, in 2011, more complex neural networks performed worse than GAMs compared to smaller neural networks. The complexity-year interactions for medium and large neural networks are positive and significant; this indicates that larger neural networks have been improving over time relative to small neural networks.
Figure 6 displays the marginal effects of the interaction between year and neural network complexity for the log RMSE ratio produced using the marginaleffects R package (Arel-Bundock et al.,, 2024). Predicted values and 95% confidence bands were calculated using the empirical distribution of all other covariates, providing a marginalised view over the observed data. This highlights how the performance of GAMs relative to neural networks changes over time across different complexity levels (where, again, positive values indicate that neural networks perform better). While performance for small and large networks remains relatively stable, medium complexity neural networks show an improvement over time. In contrast, datasets using neural networks of unknown complexity in earlier years lie just above the zero line, and this trend declines slightly over time. For all complexities in 2024, the confidence bands include zero, indicating no clear performance advantage of any neural network complexity over GAMs at present. Rug marks along the x-axis indicate the years for which there are observations for each complexity level, highlighting where data are available and model estimates are most reliable.
Table 9: Estimates from mixed-effects models.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept)  0.11 (0.28) 1.84-1.84 (1.61) 0.70-0.70 (1.11)
Year $\phantom{-}0.02$ (0.03) $\phantom{-}0.35$ (0.21) $\phantom{-}0.11$ (0.11)
[…]
Sample_Sizemedium $\phantom{-}0.33$ (0.33) $\phantom{-}3.59$ (0.27) $\phantom{-}0.52$ (0.59)
Sample_Sizelarge $\phantom{-}0.86$ (0.44) $\phantom{-}2.76$ (0.24) $\phantom{-}1.03$ (1.34)
Sample_Sizeunknown $\phantom{-}1.25$ $(0.35)^{*}$ $\phantom{-}2.55$ (0.76) $\phantom{-}0.25$ (0.86)
No_Predictorsmedium $\phantom{-}0.26$ (0.23) -0.08 (0.24)  0.34 (0.43)
No_Predictorslarge  0.35 (0.34)  2.19 (0.70)  0.50 (0.82)
No_Predictorsunknown  0.11 (0.38)  0.89 (1.04)  0.25 (0.78)
Complexitymedium -1.77 $(0.38)^{*}$ -0.62 (0.62)
Complexitylarge -1.48 $(0.51)^{*}$ -0.26 (0.91)
Complexityunknown -0.55 (0.29) -1.07 (1.02)  0.19 (0.52)
[…]
Year:Sample_Sizemedium -0.02 (0.03) -0.32 (0.02) -0.02 (0.06)
Year:Sample_Sizelarge -0.06 (0.04) -0.29 (0.05) -0.07 (0.12)
Year:Sample_Sizeunknown -0.10 $(0.03)^{*}$ -0.21 (0.09)  0.00 (0.08)
Year:No_Predictorsmedium -0.03 (0.02) -0.03 (0.06) -0.03 (0.05)
Year:No_Predictorslarge -0.03 (0.04) -0.25 (0.09) -0.06 (0.08)
Year:No_Predictorsunknown -0.01 (0.04) -0.15 (0.13)  0.00 (0.08)
Year:Complexitymedium  0.11 (0.04)  0.04 (0.07)
Year:Complexitylarge  0.10 (0.05)  0.01 (0.08)
Year:Complexityunknown  0.01 (0.03) -0.06 (0.13) -0.09 (0.05)
$\sigma^{2}_{paper}$  0.27 (0.00)  1.06 (0.00)  0.79 (0.00)
$\sigma^{2}_{residual}$  0.08 (0.00)  0.03 (0.00)  0.23 (0.00)
ICC  0.77 (0.00)  0.97 (0.00)  0.77 (0.00)
  • () indicates statistical significance at the 5% level.
Figure 6: Marginal effects of the interaction between year and neural network complexity for the log RMSE ratio. Predicted values and 95% confidence bands are derived from the mixed-effects model, averaging over the empirical distribution of other covariates. Rugs along the x-axis indicate the observations within each complexity category, highlighting where model estimates are supported by data. Arrows indicate where neural networks are superior (above zero) and where GAMs are superior (below zero).
From the model output, the datasets of unknown sample size have a positive main effect on the log RMSE ratio ($\hat{\beta}$ = 1.25, SE = 0.35) and a negative interaction with year ($\hat{\beta}$ = 0.10-0.10, SE = 0.03). Thus, at the 2011 baseline, neural networks outperform GAMs for datasets with unknown sample size relative to the reference category (small). However, the advantage significantly reduces over time; in any case, there is little practical implication here, as it is unclear how the datasets of unknown size differ from those where the sample size is reported.
The ICC is high (0.77), indicating substantial between-paper variability and supporting the appropriateness of a mixed-effects model. The addition of covariates to the model does not explain the variation across papers.
7.1.3 AUC metric analysis
The AUC analysis included 111 observations from 37 papers. No outliers were identified, and all data points were included in the analysis. For neural network complexity, the large category was not represented in the data and hence omitted from the model. Small and medium categories were merged due to the minimal observations in each, and hence small/medium was used as the reference level.
Table 9 presents the estimates from the AUC model, including all covariates and pairwise interactions with time. The intercept is not statistically significant $\hat{\beta_{0}}=-1.81$ (SE = 1.61), so that, like with RMSE, there is no overall direction to the performance results, i.e., neural networks being better than GAMs, or vice versa. However, we see that sample size and the number of predictors, along with their pairwise interactions with time, do have statistically significant effects, relative to the reference categories.
The marginal effects plot for different sample sizes through time over the empirical distribution is seen in Figure 7. The confidence bands for medium and unknown sample sizes include zero throughout, indicating no consistent difference in performance between GAMs and neural networks. In small datasets, GAMs initially outperform neural networks; however, this gap narrows through time. Conversely, neural networks had superior performance over GAMs for datasets with a large sample size, and this gap is narrowing over time. Note that neither the small nor the large datasets are statistically significant in recent years.
Figure 7: Marginal effects of the interaction between year and sample size for the log AUC loss ratio. Predicted values and 95% confidence bands are derived from the mixed-effects model, averaging over the empirical distribution of other covariates. Rugs along the x-axis indicate the observations within each complexity category, highlighting where model estimates are supported by data. Arrows indicate where neural networks are superior (above zero) and where GAMs are superior (below zero).
The interaction between the number of predictors and year on the log AUC loss ratio is shown in Figure 8. Predicted marginal effects indicate minimal temporal change across different numbers of predictors. Datasets with a medium number of predictors show the most notable change through time. GAMs outperform neural networks in earlier years, with the gap narrowing to equivalence in 2024.
Although our analysis reveals some signals for variables impacting comparative model performance, the big variability in results across different papers is particularly noteworthy; as with the RMSE analysis, the ICC is large here (ICC =0.97=0.97). Thus, results on datasets from the same paper are highly correlated with each other, indicating differences in the modelling approaches and datasets specific to individual papers that have not been quantified.
Figure 8: Marginal effects of the interaction between year and number of predictors for the log AUC loss ratio. Predicted values and 95% confidence bands are derived from the mixed-effects model, averaging over the empirical distribution of other covariates. Rugs along the x-axis indicate the observations within each complexity category, highlighting where model estimates are supported by data. Arrows indicate where neural networks are superior (above zero) and where GAMs are superior (below zero).
7.1.4 $R^{2}$ metric analysis
Datasets utilising $R^{2}$ as a performance metric totalled 141, originating from 55 papers. The log $R^{2}$ loss ratio was modelled in the same manner as the above analyses. The inclusion of predictor variables and interactions provided no evidence of driving factors on the $R^{2}$ performance as seen in Table 9.
8 Conclusion
This systematic review and performance metric analysis evaluated 430 datasets across 143 papers comparing GAMs and neural networks, offering insights into how these two modelling frameworks have been applied and assessed in the literature. The first notable findings are the key characteristics of papers carrying out these comparisons and the substantial differences in how they are reported. Although all datasets included empirical analysis of both GAMs and neural networks, important dataset-level information, such as sample size, number of predictors, and neural network complexity, was frequently unreported. Neural network complexity, in particular, was only fully reported in 35% of datasets. The lack of standardised reporting affects transparency and reproducibility.
We extracted and analysed performance outcomes across the most commonly reported metrics: RMSE, $R^{2}$, and AUC. On average, there was no evidence of a statistically significant difference between the performance of GAMs and neural networks across datasets for all three metrics based on the values of the intercepts in the models; that is to say, there is no evidence of consistent superiority of one model class over the other. This aligns with the claims of other authors, for example, Rudin, (2019) says that no single algorithm consistently outperforms others on tabular data, with differences often being small compared to tuning effects. They suggest interpretable models can perform as well as more complex ones, though publication bias often exaggerates performance gaps.
Considering covariate effects revealed potential drivers for differences in the performance of GAMs and neural networks. This highlights that, depending on the dataset structure and computing capabilities, one model may be preferred over the other on a case-by-case basis. For RMSE, model complexity and its interaction with year played a significant role. In earlier years, medium-sized neural networks were outperformed by GAMs, while neural networks of unknown size had superior performance over GAMs. Interestingly, while neural network size showed a significant association with the response, in recent years these effects have diminished and are no longer statistically distinguishable from zero for any network size.
For AUC, the pairwise interactions between year with sample size, and year with number of predictors showed that, while neural networks have tended to outperform GAMs in larger higher-dimensional datasets, this performance gap has narrowed over time; recent years show no statistically significant differences from zero for any category. This suggests that GAMs are highly competitive, especially in smaller data settings.
Our findings offer important insights for future research. First, GAMs remain a robust and interpretable choice in many applied settings and should not be dismissed simply due to the perceived superiority of more complex models. Second, in earlier years, while the performance advantage of models appeared dependent on neural network complexity and dataset dimension, in recent years, these factors no longer have an effect on performance. Researchers can weigh up the trade-off between model accuracy and interpretability. Third, the strong intra-paper correlation (high ICC values across all metrics) demonstrates the necessity of accounting for paper-level variation when comparing models. Given the large variability in model performance across different papers, it is of interest to investigate further potentially unrecorded factors at play.
In conclusion, our study provides a comprehensive assessment of the performance of GAMs versus neural networks across a diverse set of applications using tabular data. While neural networks have shown impressive performance on complex (non-tabular) datasets, GAMs continue to offer competitive performance with added interpretability and transparency on tabular datasets. Neural networks and GAMs should be considered complementary tools within the modern data-science toolkit. Choosing between them requires careful consideration of the value placed on interpretability versus predictive gains. Notably, recent research has been conducted on combining GAMs and neural networks within hybrid models (Yang et al.,, 2021; Rügamer et al.,, 2024).
Acknowledgement
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission
Appendix A Alternative model formulation
In Section 7.1.2, the log RMSE ratio (GAM/NN) is modelled as
$\log(\mathrm{RMSE}_{\mathrm{ratio}})\sim\beta_{0}+\texttt{Year}+\texttt{Area}+\dots+(\texttt{Year}\times\texttt{Complexity})+(1|\texttt{Paper\_ID}),$ (A.8)
which is equivalent to modelling
$\log(\mathrm{RMSE}_{\mathrm{GAM}})\sim\beta_{0}+\beta_{1}\log(\mathrm{RMSE}_{\mathrm{NN}})+\texttt{Year}+\texttt{Area}+\\ \dots+(\texttt{Year}\times\texttt{Complexity})+(1|\texttt{Paper\_ID}),$ (A.9)
where $\beta_{1}$ = 1. More generally, one may estimate $\beta_{1}$ from the data. In this more general modelling framework, we model $\log(\mathrm{RMSE_{GAM}})$ and include $\log(\mathrm{RMSE_{NN}})$ as a covariate along with publication year, domain area, sample size, number of predictors, and neural network complexity. Results from this model are reported in Table 10. (Models for AUC and $R^{2}$ are also shown.) The estimated coefficient for the covariate $\log(\mathrm{RMSE_{NN}})$ is $\hat{\beta_{1}}=0.99$ (SE = 0.01). There is no evidence that this estimate is statistically significantly different from 1 at the 5% level. This implies that the RMSE of GAMs and neural networks changes in a one-to-one proportion. Importantly, the set of significant predictors and their interpretations remain identical to those in Table 9.
Repeating this alternative modelling structure for AUC, we come to similar conclusions with the estimated coefficient for $\log(\mathrm{(1-AUC_{NN})})$ being $\hat{\beta_{1}}=1.03$ (SE = 0.04), which is not statistically significantly different from 1 at the 5% level. The set of significant predictors is the same as the modelling framework in the main text. For $R^{2}$, the estimate for $\log(\mathrm{(1-R^{2}_{NN})})$ is $\hat{\beta_{1}}=0.60$ (SE = 0.06), which is statistically significantly different from 1 at the 5% level. However, similar to the model in the main text, there is no evidence of any predictors having an effect on the response.
Table 10: Estimates from alternative mixed-effects model.
RMSE AUC $\boldsymbol{R}^{\boldsymbol{2}}$
Term Estimate (SE) Estimate (SE) Estimate (SE)
(Intercept)  0.11 (0.27) -1.84 (1.64) -0.71 (0.97)
$\log(\mathrm{Metric_{NN}})$  0.99 (0.01)  1.03 (0.04)  0.60 (0.06)
Year  0.02 (0.03)  0.36 (0.21) -0.02 (0.09)
[…]
Sample_Sizemedium  0.33 (0.33)  3.63 (0.27)  0.50 (0.52)
Sample_Sizelarge  0.80 (0.44)  2.81 (0.25)  0.69 (1.16)
Sample_Sizeunknown  1.28 (0.35)  2.61 (0.78)  0.44 (0.75)
No_Predictorsmedium  0.26 (0.22) -0.08 (0.24)  0.09 (0.37)
No_Predictorslarge  0.37 (0.34)  2.25 (0.72)  0.08 (0.72)
No_Predictorsunknown  0.10 (0.38)  0.88 (1.06) -0.20 (0.68)
Complexitymedium -1.74 (0.38) -0.51 (0.54)
Complexitylarge -1.46 (0.51) -0.29 (0.78)
Complexityunknown -0.54 (0.29) -1.05 (1.04)  0.08 (0.46)
[…]
Year:Sample_Sizemedium -0.02 (0.03) -0.32 (0.02) -0.01 (0.05)
Year:Sample_Sizelarge -0.06 (0.04) -0.30 (0.05) -0.04 (0.10)
Year:Sample_Sizeunknown -0.10 (0.03) -0.21 (0.09) -0.01 (0.07)
Year:No_Predictorsmedium -0.03 (0.02) -0.03 (0.06) -0.03 (0.05)
Year:No_Predictorslarge -0.04 (0.04) -0.25 (0.09)  0.00 (0.07)
Year:No_Predictorsunknown -0.01 (0.04) -0.16 (0.13)  0.06 (0.07)
Year:Complexitymedium  0.11 (0.04)  0.05 (0.06)
Year:Complexitylarge  0.10 (0.05)  0.01 (0.07)
Year:Complexityunknown  0.01 (0.03) -0.06 (0.13) -0.03 (0.05)
$\sigma^{2}_{paper}$  0.26 (0.00)  1.10 (0.00)  0.62 (0.00)
$\sigma^{2}_{residual}$  0.08 (0.00)  0.03 (0.00)  0.16 (0.00)
ICC  0.76 (0.00)  0.97 (0.00)  0.79 (0.00)
  • () indicates statistical significance at the 5% level.